# Unbiased General Annotated Dataset Generation

### v1

@@ 4.3 Robustness Against Specific Bias
In this section, we aim to figure out whether our ubGen data can help the backbones to learn good features instead of capturing specific bias as a shortcut. Hence, we follow one recent study [48] to test shape-texture bias, context bias, and background bias of the backbone networks. All these results are given in Tab. 3.
Shape-Texture Bias. Prior work shows humans primarily use shape for object recognition [49, 9], while neural networks often rely on texture cues [12, 13]. Hence, we evaluate our ubGen data’s ability to reduce texture bias using the Cue Conflict, where shape and texture cues intentionally conflict across 1200 images from 16 classes. We use $TI$ which represents the texture inclination of the model to understand the decision-making of the model when facing a shape-texture conflicting image (e.g. a cat with the texture of an elephant). Our findings indicate that training on our ubGen images, models tend to be less texture-biased. Concretely speaking, the two types of models trained on our data show $4.8\%$ and $9.8\%$ texture inclination decline compared with those trained on real images.
@@
In our evaluation result, we find the models trained on our ubGen ImageNet demonstrate leading object recognition capabilities where achieve $64.7\%$ and $66.0\%$ average relative accuracy on each backbone in a constantly changing context.
@@
IN100-ubGen $\checkmark$ $\checkmark$ $\checkmark$ Large $66.0$ $44.8$ $34.3$ $59.1$ $92.7$ $71.6$ $72.0$ $50.8$ $42.3$ $58.5$
@@
[…]
@@ 4.4 Ablation Study
In this section, we investigate the design choices of ubGen training process. Due to computational cost, without loss of generality, we conducted the ablation study on the smaller ImageNet-100 datasets [50] for evaluation. Unless otherwise specified, we choose ResNet50 backbone and mainly report the results of transfer learning.
Effect of Bi-Level Semantic Alignment Loss. According to Tab. 4, we explore the effect of the bi-level semantic alignment loss on mitigating bias. As this loss consists of two different levels of losses, we analyze the impact of each respectively. It is notable that without the entire semantic alignment loss, the average accuracy on transfer learning datasets decreased by $4.9\%$. Furthermore, when we remove individual semantic alignment loss from ubGen fine-tuning progress, the accuracy on IN100-val and transfer learning shows a significant decline. In our analysis, it appears that the absence of individual semantic alignment leads the model to solely learn the overall semantic distribution of the dataset, resulting in a lack of distinction or specific semantic meaning among different classes (see the left picture of each pair in Fig. 4). This ultimately causes a collapse when using classification as the object to train the backbone.
@@
Effect of Quality Assurance Loss. As displayed in Fig. 5, adding quality assurance loss sufficiently solves the quality deterioration when only using bi-level semantic alignment loss to optimize. Moreover, quantitative results in Tab. 4 indicate that the accuracy on both ImageNet100 and transfer learning tasks could decline in the absence of quality assurance loss.
Effect of CLIP’s Capacity. As depicted in Tab. 4, we aim to explore the effect of the knowledge of CLIP model as it plays a pivotal role in our method. Results indicate that the capacity of CLIP matters, the average accuracy decreased to $56.7\%$ when changing to a smaller size which contains lower image-text alignment capacity than a larger one.
@@ 5 Conclusion
In this study, we take the first attempt to directly generate an unbiased annotated dataset for better backbone network pre-training. In a specific, we develop a novel bi-level semantic alignment loss, which not only forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset in an adversarial learning manner, but also requires each generated image to match the semantic description of its category name. Through fine-tuning the pre-trained diffusion model with the proposed loss together with a quality assurance loss, we can obtain an unbiased annotated dataset generation model using only all category names in the target dataset as input. Experiments on different computer vision tasks demonstrate that pre-training backbone network on our generated dataset can lead to stable generalization capacity enhancement.
Appendix
Appendix A Loss Computation Algorithm
Algorithm 1 A complete loss computation step for the ubGen generator during fine-tuning
1:  $im=\text{GenerateImage}(\epsilon_{\theta},\xi,c)$
2:  $f_{te}=\text{RandomlySelect}(\{f_{c_{1}},\dots,f_{c_{1000}}\})$
3:  $f_{im},f_{p_{c}}=\text{GetFeatures}(\mathcal{C},im,p_{c})$
5:  $\mathcal{L}_{in}=\text{ComputeIndividualLoss}(f_{im},f_{p_{c}})$
6:  $\mathcal{L}_{bi}=\mathcal{L}_{en}+\mathcal{L}_{in}$
7:  $\mathcal{L}_{q},=\text{ComputeQualityLoss}(\mathcal{Q},im)$
8:  $\mathcal{L}=\mathcal{L}_{bi}+\lambda_{1}\mathcal{L}_{q}$
Output: Training loss for ubGen generator $\mathcal{L}$.
Name SD1.5
Dataset Generator
Learning rate 2e-5
Learning rate scheduler Constant
LR warmup steps 0
Optimizer AdamW
AdamW - $\beta_{1}$ 0.9
AdamW - $\beta_{2}$ 0.999
Gradient clipping $0.1$
Discriminator
Learning rate 1e-5
Learning rate scheduler Constant
Optimizer AdamW
AdamW - $\beta_{1}$ 00
AdamW - $\beta_{2}$ $0.999$
Gradient clipping $1.0$
Quality assurance loss weight $\lambda_{2}$ 0.1
Gradient enable steps $5$
LoRA rank 128
Classifier-free guidance scale 2
Resolution $512\times 512$
Total training epochs 3
Local batch size $4$
Mixed Precision FP16
Table 5: ubGen training hyperparameters for SD1.5.
Appendix B Scoring Quality
[…]
$p_{l_{i}}=\frac{e^{\mathcal{X}_{l_{i}}}}{\sum_{j=1}^{5}{e^{\mathcal{X}_{l_{j}}}}}.$ (10)
where $S_{q}$ is ranging from one to five.
@@ Appendix C Training Details
Appendix D Datasets Details
Except for ImageNet, We also compare with other two synthetic ImageNet datasets [1, 60] because they are the only open source datasets based on SD1.5. Thus, we can get fairer and more convincing results based on one implementation. In addition, all datasets used in our metrics to benchmark the bias of the datasets and test the generalization capacities of the backbones are listed in the Tab. 7.
Appendix E Data Synthesis Details
We use SD1.5 [43] across all benchmarks. Besides, text prompt “classnames" and hyperparameters showd in Tab. 8 are used to synthesize ImageNet-like datasets (IN-1k, IN-100).
@@ Appendix E Data Synthesis Details
Appendix F Limitation
While our ubGen demonstrates a great potential to obtain unbiased annotated dataset like ImageNet, the polysemy of some text descriptions may bring drawbacks. As shown in Fig. 6, some divergences occur when the class name refers to several objects . For instance, the text “crane" can denote either a bird or a machine, and when prompted with “crane" to generate a class in our dataset, two entirely different objects will appear. We consider that these divergences are caused by the multiple directions of clip text space due to the polysemy of human words and may compromise the knowledge of classification models trained on our dataset. Although we believe this issue can be solved with more specific text descriptions instead of class names, how to introduce more specific text descriptions without additional bias other than object is still unclear. We will explore it in our future works.
In addition, how to apply ubGen to generate large-scale multimodal datasets and general unannotated datasets has not been answered. Addressing this will also be a direction of our future research efforts.

### v3

@@ 4.3 Robustness Against Specific Bias
In this section, we aim to figure out whether our lbGen data can help the backbones to learn good features instead of capturing specific bias as a shortcut. Hence, we follow one recent study [48] to test shape-texture bias, context bias, and background bias of the backbone networks. All these results are given in Table 3.
Shape-Texture Bias. Prior work shows humans primarily use shape for object recognition [49, 9], while neural networks often rely on texture cues [12, 13]. Hence, we evaluate whether our data can reduce texture bias using the Cue Conflict, where shape and texture cues intentionally conflict across 1200 images from 16 classes. We use $TI$ which represents the texture inclination of the model to understand the decision-making of the model when facing a shape-texture conflicting image (e.g. a cat with the texture of an elephant). Our findings indicate that training on our lbGen images, models tend to be less texture-biased. Concretely speaking, the two types of models trained on our data show $4.8\%$ and $9.8\%$ texture inclination decline compared with those trained on real images.
@@
In our evaluation result, we find the models trained on our lbGen ImageNet demonstrate leading object recognition capabilities where achieve $64.7\%$ and $66.0\%$ average relative accuracy on each backbone in a constantly changing context.
@@
IN100-lbGen $\checkmark$ $\checkmark$ $\checkmark$ Large $66.0$ $44.8$ $34.3$ $59.1$ $92.7$ $71.6$ $72.0$ $50.8$ $42.3$ $58.5$
@@
Background Bias. The background bias of models can be used to identify if the model is using the background of the image during training to improve the classification accuracy instead of using the object itself. For the two datasets that we utilize to evaluate the background bias, the Mixed-Rand segments the foreground object in an image and switches the original background with a random background from a different class label, while the Mixed-Same partition places the segmented foreground object on a random background from the same class label. Thus, we can use $BG_{Gap}$ which measures the difference in performance between these two datasets to examine how decision-making processes can be influenced just by changing the background to a different class. As we report in Table 3, The ResNet50 and ViT-S trained on our data obtain $6.4\%$ and $6.1\%$ performance gaps, which shows lower gaps compared with training on other synthetic data and real data.
@@ 4.4 Ablation Study
In this section, we investigate the design choices of lbGen training process. Due to computational cost, without loss of generality, we conducted the ablation study on the smaller ImageNet-100 datasets [50] for evaluation. Unless otherwise specified, we choose ResNet50 backbone and mainly report the results of transfer learning.
Effect of Bi-Level Semantic Alignment Loss. According to Table 4, it is notable that without the entire semantic alignment part, the average accuracy decreased by $4.9\%$. Furthermore, when we remove individual semantic alignment part, the accuracy on IN100-val and transfer learning data shows a significant decline. In our analysis, it appears that the absence of individual semantic alignment leads the model to solely learn the overall semantic distribution of the dataset, resulting in a lack of distinction or specific semantic meaning among different classes (see the left picture of each pair in Figure 4). This ultimately causes a collapse when training the backbone. In sum, the bi-level semantic alignment loss successfully help to align the generated images into low-biased semantic space from both the entire dataset distribution level and the specific object category level.
Effect of Quality Assurance Loss. As displayed in Figure 5, adding quality assurance loss sufficiently solves the quality deterioration when only using bi-level semantic alignment loss. Moreover, results in Table 4 indicate that the accuracy on both ImageNet100 and transfer learning tasks could decline in the absence of quality assurance loss. Thus, it can be drawn that the quality assurance loss successfully help to guarantee the low-level image quality during generation.
@@
Effect of CLIP’s Capacity. As depicted in Table 4, we aim to explore the effect of the knowledge of CLIP model as it plays a pivotal role in our method. Results indicate that the capacity of CLIP matters, the average accuracy decreased to $56.7\%$ when changing to a smaller size which contains lower image-text alignment capacity than a larger one.
@@ 5 Conclusion
In this study, we take the first attempt to directly generate a low-biased annotated dataset for more generalized backbone network pre-training. Specifically, we develop a novel bi-level semantic alignment loss, which not only forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset, but also requires each generated image to match the semantic description of its category. Through fine-tuning the pre-trained diffusion model with the proposed loss together with a quality assurance loss which helps to guarantee the low-level image quality, we can obtain a low-biased annotated dataset generation model using only all category names in the target dataset as input. Experiments on various tasks demonstrate that pre-training backbone network on our generated dataset can lead to stable generalization capacity enhancement.
6 Acknowledgment
This work is supported in part by the National Natural Science Foundation of China under Grand 62372379, and Grant 62472359; in part by the Xi’an’s Key Industrial Chain Core Technology Breakthrough Project: AI Core Technology Breakthrough under Grand 23ZDCYJSGG0003-2023.
Appendix of lbGen
@@ 6 Acknowledgment
Appendix A Loss Computation Algorithm
Algorithm 1 A complete loss computation step for the lbGen generator during fine-tuning
1:  $im=\text{GenerateImage}(\epsilon_{\theta},\xi,c)$
2:  $f_{te}=\text{RandomlySelect}(\{f_{c_{1}},\dots,f_{c_{1000}}\})$
3:  $f_{im},f_{p_{c}}=\text{GetFeatures}(\mathcal{C},im,p_{c})$
5:  $\mathcal{L}_{in}=\text{ComputeIndividualLoss}(f_{im},f_{p_{c}})$
6:  $\mathcal{L}_{bi}=\mathcal{L}_{en}+\mathcal{L}_{in}$
7:  $\mathcal{L}_{q},=\text{ComputeQualityLoss}(\mathcal{Q},im)$
8:  $\mathcal{L}=\mathcal{L}_{bi}+\lambda_{1}\mathcal{L}_{q}$
Output: Training loss for lbGen generator $\mathcal{L}$.
Appendix B Scoring Quality
[…]
$p_{l_{i}}=\frac{e^{\mathcal{X}_{l_{i}}}}{\sum_{j=1}^{5}{e^{\mathcal{X}_{l_{j}}}}}.$ (10)
where $S_{q}$ is ranging from one to five.
@@ Appendix C Training Details
Appendix D Data Synthesis Details
We use SD1.5 [43] across all benchmarks. Besides, text prompt “classnames" and hyperparameters showd in Table 7 are used to synthesize ImageNet-like datasets (IN-1k, IN-100).
@@ Appendix D Data Synthesis Details
Appendix E Datasets Details
Except for ImageNet, We also compare with other two synthetic ImageNet datasets [1, 60] because they are the only open source datasets based on SD1.5. Thus, we can get fairer and more convincing results based on one implementation. In addition, all datasets used in our metrics to benchmark the bias of the datasets and test the generalization capacities of the backbones are listed in the Table 6.
Name SD1.5
Dataset Generator
Learning rate 2e-5
Learning rate scheduler Constant
LR warmup steps 0
Optimizer AdamW
AdamW - $\beta_{1}$ 0.9
AdamW - $\beta_{2}$ 0.999
Gradient clipping $0.1$
Discriminator
Learning rate 1e-5
Learning rate scheduler Constant
Optimizer AdamW
AdamW - $\beta_{1}$ 00
AdamW - $\beta_{2}$ $0.999$
Gradient clipping $1.0$
Quality assurance loss weight $\lambda_{2}$ 0.1
Gradient enable steps $5$
LoRA rank 128
Classifier-free guidance scale 2
Resolution $512\times 512$
Total training epochs 3
Local batch size $4$
Mixed Precision FP16
Table 8: lbGen training hyperparameters for SD1.5.
Appendix F Computing Resources
It takes about 1 hour to fine-tune the generator and 52 hours to generate the ImageNet-like dataset ($\sim$1.3M images) with 8 A100 GPUs. The generation runtime of each image is comparable to existing diffusion models.
Appendix G Limitation
While our lbGen demonstrates a great potential to obtain low-biased annotated dataset like ImageNet, the polysemy of some text descriptions may bring drawbacks. As shown in Figure  6, some divergences occur when the class name refers to several objects . For instance, the text “crane" can denote either a bird or a machine, and when prompted with “crane" to generate a class in our dataset, two entirely different objects will appear. We consider that these divergences are caused by the multiple directions of clip text space due to the polysemy of human words and may compromise the knowledge of classification models trained on our dataset. Although we believe this issue can be solved with more specific text descriptions instead of class names, how to introduce more specific text descriptions without additional bias other than object is still unclear. We will explore it in our future works.
@@ Appendix G Limitation
What’s more, our method attempts employing the low-biased text information (e.g., object category name) to regularize and fine-tune the diffusion model in the CLIP feature space for low-biased image generation. Although the diffusion model is only fine-tuned on the 1K categories in ImageNet, our generated dataset shows less bias (i.e., better generalization capacity in downstream tasks) than other competitors. However, on one hand, since the fine-grained categories in ImageNet are scarce, the generalization performance of our method in fine-grained object recognition tasks is still limited. On the other hand, compared with the infinite categories of objects in real world, the number of categories employed for fine-tuning remains limited. This also restrict the generalization capacity of our method, i.e., produces bias. Fortunately, our method provides a general low-biased dataset generation framework, which can mitigate both limitations mentioned above by simply introducing more object categories for fine-tuning.